# Elevating Software Trust: Unveiling and Quantifying the Risk Landscape

### v1

@@ V-A3 Programming Language
Motivation: Literature indicates that extensive experience within a domain is essential to achieve exceptional levels of performance [54]. In other words, trust and distrust towards an entity rely on the experience. The subtraction from 1 indicates the inverse effect, i.e., as the experience proportion increases, the risk associated with it decreases.
@@
To compute the developer’s overall risk ($R^{\textup{DEV}}_{S_{i},t_{k}}$), the weighted sum of three individual components, i.e., the risks related to the code dependencies ($R^{\textup{CD}}_{S_{i},t_{k}}$), code specifications ($R^{\textup{CS}}_{S_{i},t_{k}}$), and programming language ($R^{\textup{PL}}_{S_{i},t_{k}}$) have been calculated as:
@@
Motivation: Code coverage quantifies how thoroughly a test suite tests a software system [55]. Literature suggests that code coverage helps in maintenance tasks like predicting the likelihood of faults in a software [56, 57].
@@
Motivation: As mentioned above, the expertise and experience of the developer play a significant role in developing high-performing software [54]. Exponential function with a negative power is used to inversely emphasize on the smaller positive values. Moreover, in comparison to alternative functions such as $\tanh$, it maps all real numbers to positive real numbers and is simpler to scale its value between 0 and 1.
@@
$E^{\textup{DB}}_{S_{i},t_{k}}=\frac{|S^{\textup{L}}_{t_{k}}|}{|S^{\textup{D}}_{t_{k}}|},$ (8)
where $S^{\textup{L}}_{t_{k}}$ is the set of all previously developed software by the same developer in the same language, and $S^{\textup{D}}_{t_{k}}$ is the set of all previously developed software by the same developer.
@@ V-B Publisher based Risk:
Assumption: Similar to the developer’s experience, the experience and expertise of all actors involved impact the risk associated with software. The average number of software published per year and the frequency of updates required for that software are inversely related to the risk segment.
@@
$R^{\textup{PB}}_{S_{i},t_{k}}=\begin{cases}\frac{\sum_{x}E^{\textup{PY}}_{x,S_{i},t_{k}}}{\sum_{x}E^{\textup{PX}}_{x,S_{i},t_{k}}}.\frac{1}{F^{\textup{UP}}_{S_{i},t_{k}}},&\textit{if $E^{\textup{PX}}_{x,S_{i},t_{k}}>0$},\\ 1,&\textit{if $E^{\textup{PX}}_{x,S_{i},t_{k}}=0$},\end{cases}$ (10)
@@ V-C User indicated Risk:
Motivation: The observed distribution of data suggests that consumer trust in software is represented by the ratio of users who expressed satisfaction with the software to those who downloaded it. The subtraction from 1 implies the inverse effect, i.e., as the satisfaction proportion increases, the risk associated with it decreases.
@@ V-D Penalty $(0\leq P_{S_{i},t_{k}}\leq 1)$:
Motivation: Exponential expression is used to emphasize the smaller positive values. The subtraction from 1 specifies the vulnerabilities result in an inverse effect, i.e., as the exponential expression increases, the penalty associated with it decreases. It is noteworthy that for the same base value, a higher proportion of unfixed vulnerabilities results in a higher value, whereas if the base value is increased, the impact of the unresolved vulnerabilities is less pronounced. The context parameter is employed to adjust the stringency of penalties based on the software category, while the proportion of unresolved vulnerabilities reflects the level of susceptibility. Literature in the trust management domain suggests the use of pre-defined fixed values for penalty factor. However, in the proposed risk evaluation framework, we aim for a data-driven and dynamic penalty factor [59].
@@ V-E Final Risk Score $(0\leq R^{\textup{F}}_{S_{i},t_{k}}\leq 1)$:
The final risk score ($R^{\textup{F}}_{S_{i},t_{k}}$) combines the actor (i.e., developer, publisher, and user) based risk components into a single score and is computed as $w^{\textup{DEV}}_{S_{i},t_{k}}.R^{\textup{DEV}}_{S_{i},t_{k}}+w^{\textup{PB}}_{S_{i},t_{k}}.R^{\textup{PB}}_{S_{i},t_{k}}+w^{\textup{UR}}_{S_{i},t_{k}}.R^{\textup{UR}}_{S_{i},t_{k}}$. However, to enable comparisons among various risk scores, the risk has been depicted on a scale ranging from 0 to 1 by employing the sigmoid function [60]. This guarantees uniform interpretation for comparability. Moreover, to make the resulting score gradually increase, the sigmoid function has been modified, and the adjustment factors, e.g., 4 and 0.04, have been determined by graph analysis to analyze the potential outcomes. The resultant final risk $R^{\textup{F}}_{S_{i},t_{k}}$ is depicted as:
@@ VI Parameter Tuning
The envisaged SRiQT framework is adaptable to suit the particular needs of the organization. Moreover, it is essential to address specific scenarios (e.g., when there exists no dependencies or the publisher is a new publisher) that may arise. This section elaborates on both the customization options and occurrences of particular situations that may require modification to the risk assessment process. Please note that some customization options can add subjectivity between organizations; however, it will keep the computations within the organization non-subjective and consistent as the tuning will represent the context and not partiality.
@@ VII Results
In this section, we provide insights into the dataset utilized for calculating risk through the proposed SRiQT framework. Furthermore, yielded results and discussions on those results are also presented.
@@ VII-A Dataset
While it is technically feasible to extract data for these remaining attributes from GitHub and NPM repositories, the process involves complex data mining and analysis techniques beyond this current study’s scope. We have generated representative values for these parameters to demonstrate the efficacy of our proposed risk analysis framework without compromising its validity. This approach allows us to validate the framework’s functionality while maintaining focus on its core methodological contributions.
Let $\mathcal{D}$ denote the entire dataset and $\mathcal{D}_{x}\subset\mathcal{D}$ denote the subset of the dataset, excluding the parameters and their corresponding values that were generated. Please note that $\mathcal{D}_{x}$ only includes code dependencies, code coverage, user ratings, forks, vulnerabilities resolved, and total vulnerabilities. Table III provides the entire dataset’s ($\mathcal{D}$) structure and a few samples.
VII-B Validation
The most significant challenge when devising the proposed risk assessment framework is the unavailability of a labeled dataset, i.e., a dataset containing ground truth values with an emphasis on the trust perspective and the human element. It hinders the feasibility of validation through traditional ways, i.e., through comparing the results with the established ground truth. Therefore, conventional validation methods relying on labels or ground truth assessments are not applicable in our case. This emphasizes the necessity for alternative validation approaches. Accordingly, thorough testing was conducted to verify the effectiveness of the envisaged SRiQT framework, employing a set of 9000 samples. Through this rigorous testing, the framework’s performance is thoroughly examined, and its ability to handle various situations is assessed. The testing process is pivotal in confirming the strength and relevance of our results to broader contexts. Moreover, comparisons of the scores generated by the OpenSSF scorecard, OWASP risk calculator, and the proposed SRiQT framework have been provided to validate the necessity of the parameters employed by the proposed SRiQT framework.
Table IV presents the final outcomes via the proposed SRiQT framework for the selected samples. Table V exhibits the comparison of scores computed by OpenSSF scorecard [35] and the proposed risk framework along with the manual verification. Moreover, Table VI presents the comparison of the risk scores calculated using the proposed SRiQT framework by first including and then excluding the parameters for which self-generated data points were utilized, whereas Table VII delineates the risk scores computed by 5 assessors using OWASP risk calculator [36].
@@ VIII Discussion
This section explores how well the proposed SRiQT framework aligns with its stated contributions in Section I followed by a discussion on computed outputs.
@@ VIII-A Contributions
As mentioned earlier, the envisaged SRiQT framework aimed to (i) quantify risk to eliminate subjectivity and inconsistent risk assessments, (ii) explore the human element, i.e., the risk associated with the actors engaged in software supply chain, (iii) employ data-driven and dynamic weights to reflect the influence of individual parameters and metrics, and (iv) consider dependencies, when computing risk scores.
@@ VIII-A1 Quantification and Subjectivity
The proposed SRiQT framework provided a clear and measurable way to formulate and express metrics. This quantification made it easier to judge the risk severity and to prioritize risks based on severity to make objective decisions. For instance, if a user has to choose between 2 software that can achieve the same objectives, one software with a risk score of 0.3 out of 1, whereas the other with a risk score of 0.55 out of 1, it will be easier for the user to choose among the two options. This also removed inconsistencies in risk computations performed by different parties. Each individual parameter (e.g., developer’s reputation), major component (e.g., penalty), and risk segment (e.g., developer-based risk) has been represented in numeric value. All equations, i.e., equations 119 yield quantitative values.
@@ VIII-A3 Data-driven Weights
While developing the envisaged SRiQT framework, data-driven weights have been employed instead of pre-defined static weights as evident from equations 3, 6, 7, etc. The adaptability of dynamic weight allocation guarantees that the risk assessment remains relevant and reliable in the face of emerging threats and evolving operational environments by assessing a risk score based on the instantaneous values of the contributing parameters. In other words, the changes in the values of parameters used in the computation of weights associated with risk segments will also impact the influence of these risk segments in the final risk score.
@@ VIII-B1 Final Risk Score
The risk associated with the developer for Sample 1 in Table IV is assessed as low and the risk attributed to the publisher and the user for the same software reaches the moderate and the high level, respectively, primarily attributed to the publisher’s lack of experience and a significantly lower number of users rating the software. Nevertheless, the impact of this high risk is mitigated in the final risk score $R^{F}$ due to a finely proportioned resolution of vulnerabilities.
The elevated developer risk in Sample 4501 arises from a heightened code specification-related risk, accompanied by a suboptimal proportion of vulnerability resolution. The software is designed for other applications (i.e., context) and, therefore, adheres to a less stringent criteria for risk assessment.
Table V: Comparison - OpenSSF Scorecard Risk Vs. Proposed SRiQT Framework Risk.
@@
VIII-C Comparison
Table V presents 10 random samples from the results generated by utilizing the OpenSSF scorecard alongside the corresponding risk scores derived from the proposed risk assessment framework SRiQT for the same repositories. While comparing the outcomes yielded by both methods, it was observed that the scores computed by the two are not directly correlated. Some scores computed by the OpenSSF scorecard suggest a higher risk band as compared to the risk scores computed by SRiQT, whereas some recommend a lower risk band as opposed to the calculations made by using SRiQT. This indicates that the parameters or risk factors not considered by the OpenSSF scorecard have an impact on the final risk scores. This implies that the attributes taken into account by the OpenSSF scorecard to assess the security state of the software are insufficient for accurately determining the risk level associated with the software. It necessitates additional parameters to be taken into account to properly identify the risk status and make it more stringent. Please note that the bands utilized by OpenSSF scorecard for final score are low ($5\leq score\leq 10$), medium ($52.5\leq score<$ 5), high ($2.51.5\leq score<$ 2.5), critical ($score<1.5$) represented by colored rings of yellow, orange, red, and maroon respectively on the scores, whereas the risk bands applied on the final risk score in the proposed methodology are low ($<0.25$), moderate ($0.50.25\leq risk<$ 0.5), high ($0.750.5\leq risk<$ 0.75), and critical ($0.75\leq risk\leq 1$). Sample 3 has been scored as a medium risk by the OpenSSF scorecard (column 3) despite having no vulnerabilities, no dangerous workflow patterns, no defined security policy, etc. Similarly, Sample 9 has been scored as low risk by the OpenSSF scorecard even though vulnerabilities exist, the code review has less than 50% changes approved, no fuzzing or static code analysis has been used, and there is no best practices badge associated. Proper formulation instead of pre-defined values for assigning scores and severity weight to individual parameters is equally as important as selecting the most relevant influencing parameters, the developer’s and publisher’s reputation and experience, along with user experience/satisfaction.
Please note that the manual verification here includes all the parameters considered by the proposed risk framework, along with the the parameters employed by the OpenSSF scorecard. Sample 1 has the following: (1) no code coverage, (2) the developer has a reputation for developing software with a significantly high number of vulnerabilities, (3) the developer has little experience in developing software using this language, (4) no fuzzing, (5) no security policy, (6) no best practices, and (7) no information on dangerous workflows, (8) high number of detected vulnerabilities, (9) no branch protection, etc. Therefore, manual verification suggests that the risk level is critical.
Similarly, Sample 8 has the following: (1) no code coverage, (2) the developer has a reputation of developing software with moderate number of vulnerabilities, (3) the developer has little experience in the developing software using this language, (4) the context of the software is other than security or automation, (5) no fuzzing, (6) no security policy, (7) no best practices, and (8) no information on dangerous workflows etc. Accordingly, manual verification suggests that the risk level is high. Please note that the manual verification risk rating is not according to the number of quoted risk factors. It is according to the severity of the risk factors and the scoring against them.

### v2

@@ V-A3 Programming Language
Motivation: Literature indicates that extensive experience within a domain is essential to achieve exceptional levels of performance [56]. In other words, trust and distrust towards an entity rely on the experience. The property of substitutability has been employed and the subtraction from 1 indicates the inverse effect, i.e., as the experience proportion increases, the risk associated with it decreases.
@@
To compute the developer’s overall risk ($R^{\textup{DEV}}_{S_{i},t_{k}}$), the individual risk segments i.e., the risks related to the code dependencies ($R^{\textup{CD}}_{S_{i},t_{k}}$), code specifications ($R^{\textup{CS}}_{S_{i},t_{k}}$), and programming language ($R^{\textup{PL}}_{S_{i},t_{k}}$) have been aggregated by applying logic neutrality in LSP process, i.e., the weighted sum of individual risk segments as:
@@
Motivation: Code coverage quantifies how thoroughly a test suite tests a software system [57]. Literature suggests that code coverage helps in maintenance tasks like predicting the likelihood of faults in a software [58, 59]. The subtraction from 1 indicates the inverse effect, i.e., as the code coverage increases, the impact of the associated risk decreases.
@@
Motivation: As mentioned above, the expertise and experience of the developer play a significant role in developing high-performing software [56]. An exponential function with a negative power is used to emphasize the smaller positive values inversely. Moreover, compared to alternative functions such as $\tanh$, it maps all real numbers to positive real numbers and is more straightforward to scale its value between 0 and 1.
@@
$E^{\textup{DB}}_{S_{i},t_{k}}=\frac{\sum_{j}|S^{\textup{L}}_{j,t_{k}}|}{\sum_{j}|S^{\textup{D}}_{j,t_{k}}|},$ (8)
where $S^{\textup{L}}_{t_{k}}$ is the set of all previously developed software by the same developer in the same language, $j$ denotes a developer, and $S^{\textup{D}}_{t_{k}}$ is the set of all previously developed software by the same developer.
@@ V-B Publisher based Risk:
Assumption: Similar to the developer’s experience, the experience and expertise of all actors involved impact the risk associated with software. The average number of software published per year and the frequency of updates required for that software are inversely related to the risk segment. The property of substitutability has been utilized while applying the LSP process.
@@
$R^{\textup{PB}}_{S_{i},t_{k}}=\begin{cases}\frac{\sum_{x}E^{\textup{PY}}_{x,S_{i},t_{k}}}{\sum_{x}E^{\textup{PX}}_{x,S_{i},t_{k}}}.\frac{1}{F^{\textup{UP}}_{S_{i},t_{k}}},&\textit{if $E^{\textup{PX}}_{x,S_{i},t_{k}},F^{\textup{UP}}_{S_{i},t_{k}}\neq 0$},\\ 1,&\textit{if $E^{\textup{PX}}_{x,S_{i},t_{k}},F^{\textup{UP}}_{S_{i},t_{k}}=0$},\end{cases}$ (10)
@@ V-C User indicated Risk:
Motivation: The observed distribution of data suggests that consumer trust in software is represented by the ratio of users who expressed satisfaction with the software to those who downloaded it. The property of substitutability has been applied, and the subtraction from 1 implies the inverse effect, i.e., as the satisfaction proportion increases, the risk associated with it decreases.
@@ V-D Penalty $(0\leq P_{S_{i},t_{k}}\leq 1)$:
Motivation: Exponential expression is used to emphasize the smaller positive values. The subtraction from 1 specifies the vulnerabilities result in an inverse effect, i.e., as the exponential expression increases, the penalty associated with it decreases. It is noteworthy that for the same base value, a higher proportion of unfixed vulnerabilities results in a higher value, whereas if the base value is increased, the impact of the unresolved vulnerabilities is less pronounced. The context parameter is employed to adjust the stringency of penalties based on the software category, while the proportion of unresolved vulnerabilities reflects the level of susceptibility. Literature in the trust management domain suggests the use of pre-defined fixed values for penalty factors. However, in the proposed risk evaluation framework, we aim for a data-driven and dynamic penalty factor [61]. Context is taken as a mandatory input attribute to compute penalty while applying the LSP process. However, if the context is unknown, it is assumed that the software is security software for the most rigorous risk assessment.
@@ V-E Final Risk Score $(0\leq R^{\textup{F}}_{S_{i},t_{k}}\leq 1)$:
The final risk score ($R^{\textup{F}}_{S_{i},t_{k}}$) combines the actor (i.e., developer, publisher, and user) based risk components into a single score by applying logic neutrality in LSP process, i.e., the weighted sum of individual actor-based risk scores as $w^{\textup{DEV}}_{S_{i},t_{k}}.R^{\textup{DEV}}_{S_{i},t_{k}}+w^{\textup{PB}}_{S_{i},t_{k}}.R^{\textup{PB}}_{S_{i},t_{k}}+w^{\textup{UR}}_{S_{i},t_{k}}.R^{\textup{UR}}_{S_{i},t_{k}}$. However, to enable comparisons among various risk scores, the risk has been depicted on a scale ranging from 0 to 1 by employing the sigmoid function [62]. This guarantees uniform interpretation for comparability. Moreover, to make the resulting score gradually increase, the sigmoid function has been modified, and the adjustment factors, e.g., 4 and 0.04, have been determined by graph analysis to analyze the potential outcomes. The resultant final risk $R^{\textup{F}}_{S_{i},t_{k}}$ is depicted as:
@@ VI Parameter Tuning
The envisaged SAFER framework is adaptable to suit the particular needs of the organization. Moreover, it is essential to address specific scenarios (e.g., when there exists no dependencies or the publisher is a new publisher) that may arise. This section elaborates on both the customization options and occurrences of particular situations that may require modification to the risk assessment process. Please note that some customization options can add subjectivity between organizations; however, it will keep the computations within the organization non-subjective and consistent as the tuning will represent the context and not partiality.
@@ VII Results
In this section, we provide insights into the dataset utilized for calculating risk through the proposed SAFER framework. Furthermore, yielded results are also presented.
@@ VII-A Dataset
While it is technically feasible to extract data for these remaining attributes from GitHub and NPM repositories, the process involves complex data mining and analysis techniques beyond this current study’s scope. We have generated representative values for these parameters to demonstrate the efficacy of our proposed risk analysis framework without compromising its validity. This approach allows us to validate the framework’s functionality while maintaining focus on its core methodological contributions. The criterion used to generate this synthetic data has been provided in the appendix.
Let $\mathcal{D}$ denote the entire dataset and $\mathcal{D}_{x}\subset\mathcal{D}$ denote the subset of the dataset, excluding the parameters and their corresponding values that were generated. Please note that $\mathcal{D}_{x}$ only includes code dependencies, code coverage, user ratings, forks, vulnerabilities resolved, and total vulnerabilities. Table III provides the entire dataset’s ($\mathcal{D}$) structure and a few samples. More details can be found in the Appendix.
VII-B Risk Assessment
Table IV presents the final outcomes via the proposed SAFER framework for the selected samples. An example of how SAFER will be used in practice has been provided in the Appendix.
VII-C Evaluation & Comparison
The evaluation of the proposed SAFER framework has been performed based on user experience and usability tests. It is a well-established practice to conduct software evaluations through user experience. This approach ensures that the software meets usability standards and aligns with user needs and expectations [65, 66, 67, 68, 69]. The user experience has been assessed via usability, satisfaction, efficiency, learnability, and accessibility. To remove subjectivity from the opinions, a binary scale has been used for each user experience attribute, i.e., 0 represents bad user experience, whereas 1 represents good user experience.
Comparisons for the user experience, as well as risk scores generated by the OpenSSF scorecard, OWASP risk calculator, and the proposed SAFER framework, have been provided to evaluate and validate the necessity of the parameters employed by the proposed SAFER framework. Moreover, tests related to errors and missing information against parameters have been performed, and assumptions are embedded in the functionality of the framework for added convenience. More details on these tests can be found in the Appendix.
Table V outlines the comparison of user experiences associated with the OpenSSF scorecard, OWASP risk calculator, and the proposed SAFER framework for risk assessment of private repositories. Table VI exhibits the comparison of scores computed by OpenSSF scorecard [35] and the proposed SAFER framework along with the manual verification. Moreover, Figures 2 and 3 demonstrate the sensitivity and the dynamic aspect of the proposed SAFER framework in the form of heat maps, respectively. Furthermore, Table VII illustrates the comparison of the risk scores calculated using the proposed SAFER framework by first including and then excluding the parameters for which self-generated data points were utilized, whereas Table VIII delineates the risk scores computed by 5 assessors using OWASP risk calculator [36].
@@ VIII Discussion
This section explores how well the proposed SAFER framework aligns with its stated contributions in Section I followed by a discussion on computed outputs.
@@ VIII-A Contributions
As mentioned earlier, the envisaged SAFER framework aimed to (i) quantify risk to eliminate subjectivity and inconsistent risk assessments, (ii) explore the human element, i.e., the risk associated with the actors engaged in the software supply chain, (iii) employ data-driven and dynamic weights to reflect the influence of individual parameters and metrics, and (iv) consider dependencies, when computing risk scores.
@@ VIII-A1 Quantification and Subjectivity
The proposed SAFER framework provided a clear and measurable way to formulate and express metrics. This quantification made it easier to judge the risk severity and to prioritize risks based on severity to make objective decisions. For instance, if a user has to choose between 2 software that can achieve the same objectives, one software with a risk score of 0.3 out of 1, whereas the other with a risk score of 0.55 out of 1, it will be easier for the user to choose among the two options. This also removed inconsistencies in risk computations performed by different parties. Each individual parameter (e.g., developer’s reputation), major component (e.g., penalty), and risk segment (e.g., developer-based risk) has been represented in numeric value. All equations, i.e., equations 119 yield quantitative values.
@@ VIII-A3 Data-driven Weights
While developing the envisaged SAFER framework, data-driven weights have been employed instead of pre-defined static weights as evident from equations 3, 6, 7, etc. The adaptability of dynamic weight allocation guarantees that the risk assessment remains relevant and reliable in the face of emerging threats and evolving operational environments by assessing a risk score based on the instantaneous values of the contributing parameters. In other words, the changes in the values of parameters used in the computation of weights associated with risk segments will also impact the influence of these risk segments in the final risk score. For instance, $R^{CD}$ is the risk based on code dependencies, and the weight associated with it is dependent on the code coverage which implies that if the code has been tested thoroughly (e.g., code coverage $C^{C}$ is 0.96), the weightage of the risk related to code dependencies $R^{CD}$ in the $R^{DEV}$ reduces (e.g., $w^{CD}$ is 0.04) and vice versa (e.g., $w^{CD}$ is 1 when code coverage $C^{C}$ is 0). Similarly, the higher the possibility of a vulnerability remaining unresolved, the greater the influence $w^{PB}$ of the publisher-based risk $R^{PB}$ in the $R^{F}$ and vice versa.
@@ VIII-B Risk Assessments
Table V: User Experience Comparison.
Frameworks Usability Satisfaction Efficiency Learnability Accessibility
OpenSSF Scorecard 20% 100% 100% 20% 20%
OWASP Risk Calculator 20% 20% 100% 20% 100%
Proposed SAFER 100% 100% 100% 80% 100%
Figure 2: Heat Maps - Sensitivity of the Proposed SAFER Framework (a) Effect of Individual Risk Segments on Final Risk Score, (b) Residual Mapping for Effect of Individual Risk Segments on Final Risk Score, (c) Effect of Individual Parameters on Developer based Risk, and (d) Effect of Changing Values of Developer based Risk Weightage on Final Risk Score.
Figure 3: Heat Maps - Dynamicity of the Proposed SAFER Framework (a) Effect of Changing Values of weight $w^{CD}$ associated with Code Dependencies based Risk on Contribution towards $R^{DEV}$, (b) Effect of Changing Values of Publisher based Risk $R^{PB}$ on Contribution towards $R^{F}$ where column 2 and column 3 represent the values of $w^{PB}*R^{PB}$, (c) Effect of Changing Context on Penalty, and (d) Effect of Changing Proportion of Unresolved Vulnerabilities on Penalty.
@@ VIII-B1 Final Risk Score
The risk associated with the developer for Sample 1 in Table IV is assessed as low, and the risk attributed to the publisher and the user for the same software reaches the moderate and the high level, respectively, primarily attributed to the publisher’s lack of experience and a significantly lower number of users rating the software. Nevertheless, the impact of this high risk is mitigated in the final risk score $R^{F}$ due to a finely proportioned resolution of vulnerabilities.
The elevated developer risk in Sample 4501 arises from a heightened code specification-related risk, accompanied by a suboptimal proportion of vulnerability resolution. The software is designed for other applications (i.e., context) and, therefore, adheres to less stringent criteria for risk assessment.
Table VI: Comparison - OpenSSF Scorecard Risk Vs. Proposed SAFER Framework Risk.
@@
VIII-C Evaluation & Comparisons
Table V outlines the comparison of user experiences associated with the OpenSSF scorecard, OWASP risk calculator, and the proposed SAFER framework for risk assessment of private repositories. Usability refers to how easy and user-friendly the tool is to operate. Proposed SAFER: 100% of the participants rated it as easy and user-friendly to operate. OWASP: non-expert users found selecting options via the dropdown menu to be quite easy, however, due to the lack of knowledge regarding the threat agents, vulnerability factors, and/or impact factors, struggled to operate the tool effectively. OpenSSF scorecard: none of the participants had GitHub advanced security license, therefore, they needed to operate it through the command line. Most participants, specially non-expert and non-technical participants found the installation process complicated and limited support for windows’ platform posed an additional challenge. Satisfaction relates to how satisfied users are with the tool’s performance and functionality. Proposed SAFER: 100% of the participants rated it as satisfactory. OWASP: non-expert and non-technical participants were all able to compute risk scores using OWASP, however, they were not confident in the options they selected on individual dropdown menus. OpenSSF scorecard: As most of the participants couldn’t get past the installation stage, they tried OpenSSF scorecard for public repositories and 100% of the participants rated it as satisfactory. Efficiency indicates how quickly and effectively users can complete their tasks using the tool. Proposed SAFER: 100% of the participants rated it as efficient. OWASP: 100% of the participants rated it as efficient. OpenSSF scorecard: as mentioned before, majority of the participants were unable to use OpenSSF for private repositories, they rated the efficiency of OpenSSF scorecard by trying OpenSSF scorecard for public repositories and 100% of the participants rated it as efficient. Learnability describes how easy it is for new users to learn how to operate the tool. Proposed SAFER: 80% of the participants rated it as effortless. OWASP: non-expert and non-technical users found selecting options through the dropdown menu to be straightforward, however, found it challenging to understand the individual influencing parameters. OpenSSF scorecard: Majority of the participants were unable to even complete the installation stage, therefore, they couldn’t start using OpenSSF scorecard for private repositories. Accessibility refers to how accessible the tool is to users with different abilities and needs. Proposed SAFER: 100% of the participants rated it accessible. OWASP: 100% of the participants rated it accessible. OpenSSF scorecard: all the participants were able to find all the files needed for the installation process, however, the OpenSSF scorecard for private repositories itself was inaccessible for them as they couldn’t finish installation.
Figure 2 demonstrates the sensitivity of the proposed SAFER framework for risk assessment. Figure 2(a) and (b) illustrate the effect of individual risk segments, i.e., $R^{DEV}$, $R^{PB}$, and $R^{UR}$ on the final risk score $R^{FP}$ via actual $R^{FP}$ value heat map and a residual heat map, respectively. It is evident that the risk segment associated with the development phase, i.e., $R^{DEV}$, has the most significant impact on determining the final risk value. This suggests that factors and uncertainties within the development stage play a crucial role in shaping the overall outcome, making it a critical area of focus for risk management and mitigation efforts. Similarly, Figure 2(c) presents the effect of the number of vulnerabilities in all the software developed by the same developer $V^{D}$, code coverage $C^{C}$, and experience of the developer in terms of programming language $E^{DB}$ on the developer’s based risk segment $R^{DEV}$, whereas Figure 2(d) depicts the affect of the weight $w^{DEV}$ associated with the developer’s based risk segment $R^{DEV}$ on the final risk score $R^{FP}$. It can be seen that the vulnerabilities in the developed software ($V^{D}$) has the most impact on the developer’s based risk segment $R^{DEV}$ followed by the developer’s expertise in the programming language ($E^{DB}$). It can also be observed that the increasing weight $w^{DEV}$ associated with the developer’s based risk segment has an increasing impact on the final risk score $R^{FP}$.
Figure 3 exhibits the dynamic aspect of the proposed SAFER framework for risk assessment. Figure 3(a) indicates the effect of varying values of weight $w^{CD}$ associated with code dependencies based risk on contribution towards $R^{DEV}$. It can be seen that the increasing code coverage has a decreasing impact on the weight associated with $R^{CD}$, consequently reducing the contribution of $R^{CD}$ towards $R^{DEV}$. Similarly, Figure 3(b) shows the impact of changing values of publisher-based risk $R^{PB}$ on contribution towards $R^{F}$. It is apparent that a reduced publisher-based risk while keeping the weight associated with it constant has a lower contribution towards the final score $R^{F}$ and vice versa. Figure 3(c) depicts the influence of distinct context on penalty, whereas Figure 3(d) presents the effect of changing the proportion of unresolved vulnerabilities on penalty. It can be noted that the penalty values applied on the security software (i.e., $C^{TXT}$ = 0.2) are much higher than the penalty on other software while keeping the proportion of unresolved vulnerabilities ($V^{UP}$) constant. In contrast, the increasing proportion of unresolved vulnerabilities ($V^{UP}$) while keeping the context constant has an increasing impact on the penalty.